Please help with solving these problems they are measurement system conversion: I have tried using the internet and their answers are wrong. I don't know what I am doing wrong. These are questions for a Pharmacy major or Pharmacist

8.9 pounds (apothecary) = _____ drams

13 ounces = ______ grains

18 fluid drams = ______ ml

183 drams = ______ pounds (apothecary)
Answer #1

1.

Apothecary is the system used widely in the US measuring methods that introduces the unit of drams. The given equation can be solved as following-

1 pound = 116.66 drams

8.9 pounds = 116.66 * 8.9

therefore, the answer is 8.9 pounds equals to 1038.33 drams.

2.

ANother unit of mass measurement in apothecary system is ounce and grains, the converstion of following equation can be as follows-

1 ounce = 437.5 grains

therefore, 13 ounce = 437.5 * 13 = 5687.5 grains.

3.

1 fluid dram = 3.69669 ml

therefore, 18 fluid dram = 3.69669 * 18 = 66.54 ml.

4.

1 dram = 0.0016 pounds

therefore, 183 dram = 0.0016 * 183 = 0.2928 pounds.